# Gate-defined flat-band charge carrier confinement in twisted bilayer graphene

### v1

@@ Abstract
Twisted bilayer graphene (tBLG) near the magic angle is an interesting platform to study correlated electronic phases. These phases are gate-tunable and are closely related to the presence of flat electronic bands, isolated by single-particle band gaps. This allows electrostatically controlled confinement of charge carriers in the flat bands to explore the interplay between confinement, band renormalisation, electron-electron interactions and the moiré superlattice, potentially revealing key mechanisms underlying these electronic phases. Here, we show gate-controlled flat-band charge carrier confinement in near-magic-angle tBLG, resulting in well-tunable Coulomb blockade resonances arising from the charging of electrostatically defined islands in tBLG. Coulomb resonance measurements allow to study magnetic field-induced quantum oscillations in the density of states of the source-drain reservoirs, providing insight into the gate-tunable Fermi surfaces of tBLG. Comparison with tight-binding calculations emphasises the importance of displacement-field-induced band renormalisation, which is crucial for future advanced gate-tunable quantum devices and circuits in tBLG.
@@
Gate-controlled confinement of charge carriers in an effective two-dimensional (2D) electron system or in true 2D materials, such as Bernal-stacked bilayer graphene, enables the realisation of tunable quantum point contacts [1, 2], single electron transistors [3], quantum dots [4, 5, 6, 7] and related quantum devices. Such systems permit the investigation of band structure properties, the study of electron-electron interactions in the few-electron regime [8, 9], as well as probing the density of states in nearby reservoirs [10]. All this makes the realisation of gate-controlled charge confinement in twisted bilayer graphene interesting and promises additional avenues for studying its electronic properties [11]. Twisted bilayer graphene (tBLG) with a twist angle near $1^{\circ}$ not only has a rich phase diagram, including superconductivity [12, 13, 14, 15, 16, 17], strange metal phases and correlated Chern insulators [18, 19, 20, 21, 22], but also offers sizeable single-particle band gaps at finite energy, making charge carrier confinement feasible. The gate-tunability of these phases allowed the demonstration of built-in Josephson junctions [23, 12, 14], superconducting quantum interference devices [24], Aharonov-Bohm rings [25] and superconducting channels [26]. However, well-controlled charge confinements and quantum dots - all of which require more complex gating structures - have not yet been demonstrated.
Here, we show electrostatically controlled charge confinement in tBLG based on an advanced gating structure. By using a first set of top gates, we create a narrow conducting channel in the tBLG by exploiting the single-particle band gaps (on the order of several tens of meV) separating the flat bands from the remote conduction and valence band. An additional set of narrow gates across the channel allows local control of the band edge profile and enables the formation of isolated charge islands. This makes it possible to demonstrate sequential tunneling through flat-band charge islands and to perform magnetic field spectroscopy measurements that provide information about the Fermi surfaces in the nearby reservoirs and their sensitive dependence on out-of-plane electric fields, confirmed by tight-binding calculations.
@@
Figure 2: Formation of a gate-defined charge island. (a) Schematic illustration of the band edge profile along the gate-defined channel showing the formation of charge confinement between LG and RG by using the gap $\Delta_{+}$. (b) Large-range charge-stability diagram showing the conductance as a function of LG and RG measured at $V_{\mathrm{BG}}=3.5$ V and $V_{\mathrm{SG}}=3.5$ V (rhombus in Fig. 1d) with $V_{\mathrm{SD}}=100\,\mathrm{\mu V}$. (c) Line-cut showing the onset of Coulomb oscillations along LG for fixed $V_{\mathrm{RG}}=3.12$ V. (d) Zoom-in into the area marked by the black rectangle in panel (b) showing a set of diagonal Coulomb resonances indicating single-electron transport through a charge island located in between the LG and RG. (e) Schematic representation of the charge island which is tunnel-coupled to the source (S) and drain (D) leads and to the relevant gates. (f) Finite bias spectroscopy measurements for fixed $V_{\mathrm{RG}}=3.3301$ V [dashed line in panel (d)].
Finally, a gate-controlled charge island can be defined using both FGs to modulate the band edge profile along the conductive channel as illustrated in Fig. 2a. Fig. 2b shows a $V_{\mathrm{LG}}$-$V_{\mathrm{RG}}$ conductance map where we observe the transition from high conductance values (regime I, where the FGs are tuned such that we have flat-band carrier transport) to strongly suppressed conductance (regime II, where the FGs push the Fermi level into the gap $\Delta_{+}$). Focusing on a small gate voltage range applied to the LG and RG (black square in Fig. 2b and close-up in Fig. 2d), we observe well-defined conductance oscillations (see also Fig. 2c) that can be tuned equally well by the LG and RG (indicated by the slope of $\approx-1.08$ of the resonances in the $V_{\mathrm{LG}}$-$V_{\mathrm{RG}}$ plane). These periodic conductance oscillations (with a constant spacing of $\Delta V_{\mathrm{LG}}\approx 7.9\,$ mV), which we identify as Coulomb resonances, are a hallmark of the presence of tunneling barriers below both FGs, which consequently electrostatically define a charge island between the left and right FG. The occupation of this charge island, located in the flat bands, can be tuned by the capacitively coupled gates as illustrated in Fig. 2e and confirmed by finite bias spectroscopy measurements revealing characteristic Coulomb diamonds (Fig. 2f).
@@
Figure 3: Fully gate-tunable charge carrier confinement in the hole doped regime (a) Schematic illustration showing the formation of a charge island between LG and RG by using the gap $\Delta_{-}$. (b) Charge-stability diagrams showing diagonal Coulomb resonances for different combinations of BG and SG voltages highlighting the flexibility of the device. Measurements are taken at (i) $V_{\mathrm{BG}}=-3\,\mathrm{V}$ and $V_{\mathrm{SG}}=-3.5\,\mathrm{V}$ at $V_{\mathrm{SD}}=150\,\mu\mathrm{V}$. (ii) $V_{\mathrm{BG}}=-4\,\mathrm{V}$ and $V_{\mathrm{SG}}=-3.23\,\mathrm{V}$ at $V_{\mathrm{SD}}=100\,\mu\mathrm{V}$. (iii) $V_{\mathrm{SG}}=-5\,\mathrm{V}$ and (slightly asymmetric) split gate voltages of $V_{\mathrm{SG1}}=-2.14\,\mathrm{V}$ and $V_{\mathrm{SG2}}=-1.94\,\mathrm{V}$ at $V_{\mathrm{SD}}=100\,\mu\mathrm{V}$ (circle in Fig. 1f) and (iv) $V_{\mathrm{BG}}=-6\,\mathrm{V}$ and $V_{\mathrm{SG}}=-1.7\,\mathrm{V}$ at $V_{\mathrm{SD}}=100\,\mu\mathrm{V}$. (c) Line-cut along the RG extracted from the left most panel in (b) for $V_{\mathrm{BG}}=-3\,\mathrm{V}$ at fixed $V_{\mathrm{LG}}=-4.925$ V showing periodic Coulomb oscillations with an approximate peak spacing of $\Delta V_{\mathrm{RG}}\approx 7.7$ mV. (d) Lower panel: Finite bias spectroscopy measurements as function of $V_{\mathrm{LG}}$ at fixed $V_{\mathrm{BG}}=-5\,\mathrm{V}$ and $V_{\mathrm{RG}}=-3.501$ V (see dashed line in central right panel in (b)). Upper panel: Zero bias Coulomb peak trace extracted from the Coulomb diamond measurement. Note the fact that complete current suppression can be achieved in this regime.
@@
The robustness of the gate-defined charge confinement can be demonstrated by inverting all gate voltages and using the gap $\Delta_{-}$ to form local tunneling barriers, as illustrated in Fig. 3a. In Fig. 3b, we show corresponding $V_{\mathrm{LG}}$-$V_{\mathrm{RG}}$ conductance maps for four different $V_{\rm BG}$ values, where $V_{\rm SG}$ is set to ensure that the Fermi level is well within the gap $\Delta_{-}$ in the tBLG region below the SGs, and the LG and RG are fine-tuned to form tunneling barriers below the FGs (see also Supporting Section IV). In each case, we observe periodic Coulomb resonances (diagonal features), demonstrating the versatility of the device to create confinements using both single-particle band gaps of the tBLG electron system. Interestingly, we observe for all investigated regimes a nearly constant spacing of the Coulomb resonances of $\Delta V_{\mathrm{LG}}$ in the range of $(10.6\pm 3.2)$ mV (see e.g. conductance trace shown in Fig. 2c) and $\Delta V_{\mathrm{RG}}$ in the range of $(9.5\pm 2)$ mV, both consistent with the value observed for the electron-doped regime. This, in conjunction with the well-developed diagonal slopes with values in the range of $\approx-0.6$ to $\approx-1.1$ (Supporting Section II) and the possibility to completely suppress the conductance in certain gate voltage regimes, which results in the formation of well-developed Coulomb peaks (Fig. 3d), provides compelling evidence that the island is indeed formed consistently in the channel in between the two FGs. From the characteristic spacing of the Coulomb oscillations $\Delta V_{\mathrm{FG}}$ we can estimate the capacitance between the island and the FGs given by $C_{\mathrm{FG}}=e/\Delta V_{\mathrm{FG}}$ [27] resulting in $C_{\mathrm{FG}}$ in the range of $12$ to $22$ $\mathrm{aF}$. Interestingly, this value is more than a magnitude smaller than the self-capacitance extracted from the charging energy of the islands, which is consistently found to be in the range of $\approx 0.5$ meV (white dashed lines in Fig. 2f and lower panel in Fig. 3d), resulting in a self-capacitance of $C=e^{2}/E_{\mathrm{c}}\approx 320\,\mathrm{aF}$.
@@
Figure 5: Gate-tunable Fermi surface. (a) $B$-field derivative of the bulk conductance as a function of $V_{\mathrm{BG}}$ (bottom axis) or filling factor $\nu$ (top axis) and out-of-plane magnetic field. The dashed lines highlight the Landau levels emerging from CI at $\nu=-2$ (black) and the CNP at $\nu=0$ (yellow). (b) FFT of the $B$-field dependent quantum oscillations extracted for the different BG voltages (see Fig. 4d). (c) Quantum oscillation frequency $F_{B}$ obtained from tight-binding calculations for displacement field strengths of $D/\varepsilon_{0}=0$ V/nm (orange data) and $D/\varepsilon_{0}=0.48$ V/nm (dark red data). Note, that for converting $\nu$ to $V_{\mathrm{BG}}$ we neglect here stray fields from other surrounding gates. (d) Contour plots of the Fermi surface for different filling factors (i.e. energies) around the $K_{\mathrm{M}}$ points without (top row) and with (bottom row) an applied displacement field. (e) Schematic representation of the $D$-field induced energy shift. For more information see text.
@@
In conclusion, our measurements on gate-controlled electrostatic charge confinement in tBLG allow to gain insight into band deformations of (near) magic-angle tBLG in the presence of out-of-plane electric fields. Essentially, our results fill the gap between the large-angle regime ($\theta\gg 1^{\circ}$), which is characterized by large bandwidths and comparatively little sensitivity to external electrical fields [42, 23] and the small-angle regime ($\theta\ll 1^{\circ}$), where the presence of electric fields causes the formation of topological helical networks [49, 50, 51]. An interesting observation is the absence of the correlated insulating state in Fig. 5b, while it is prominently present in the bulk transport data shown in Fig. 5a.
Given the recent theoretical and experimental study on the importance of long-ranged Coulomb interactions in the formation of correlated states in (near) magic-angle tBLG [52, 16, 44], we emphasize the flexibility of our device architecture to study electric field-induced screening of electronic interactions. In that respect, future experiments closer to the magic angle may provide interesting insights into a number of correlated phenomena including superconductivity and strange metal phases. In addition, the presented technology may allow the development of complex quantum devices and circuits based on tBLG, ranging from coupled quantum dots, chains of Josephson junction-based SQUIDs to single-photon detector arrays.
@@
Author contributions C.S. and A.R. conceived the experiment. A.R. built the device and performed the measurements. A.F. performed the theoretical simulations. A.R. and A.A. performed the data analysis under the supervision of C.S.. E.I., K.H. and L.B. provided assistance with the measurement setup. M.O. performed the ALD step. S.T. and F.L. performed the electron beam lithography. K.W. and T.T. supplied the hBN crystals. A.R., A.F., R.J.D., D.M.K., B.B. and C.S. discussed the experimental and theoretical data and wrote the manuscript.
Data availability The data supporting the findings of this study are available in a Zenodo repository under https://doi.org/10.5281/zenodo.13737631.

### v2

# Gate-defined single-electron transistors in twisted bilayer graphene

@@ Abstract
Twisted bilayer graphene (tBLG) near the magic angle is a unique platform where the combination of topology and strong correlations gives rise to exotic electronic phases. These phases are gate-tunable and related to the presence of flat electronic bands, isolated by single-particle band gaps. This enables gate-controlled charge confinement, essential for the operation of single-electron transistors (SETs), and allows to explore the interplay of confinement, electron interactions, band renormalisation and the moiré superlattice, potentially revealing key paradigms of strong correlations. Here, we present gate-defined SETs in near-magic-angle tBLG with well-tunable Coulomb blockade resonances. These SETs allow to study magnetic field-induced quantum oscillations in the density of states of the source-drain reservoirs, providing insight into gate-tunable Fermi surfaces of tBLG. Comparison with tight-binding calculations highlights the importance of displacement-field-induced band renormalisation crucial for future advanced gate-tunable quantum devices and circuits in tBLG including e.g. quantum dots and Josephson junction arrays.
@@
Single-electron transistors (SETs) [1] are highly sensitive quantum electronic devices that exploit the control of individual charge carriers flowing through a small conducting island or a quantum dot. SETs have been realized in many different material systems (including van der Waals materials such as graphene [2] and 2D semiconductors [3]) and in particular with integrated quantum dots they permit the investigation of band structure properties, the study of electron-electron interactions [4, 5], as well as probing the density of states in nearby reservoirs [6]. All this makes the realisation of tunable SETs in twisted bilayer graphene interesting and promises additional avenues for studying its electronic properties. Twisted bilayer graphene (tBLG) with a twist angle near $1^{\circ}$ not only has a rich phase diagram, including superconductivity [7, 8, 9, 10, 11, 12], strange metal phases and correlated Chern insulators [13, 14, 15, 16, 17], but also offers sizeable single-particle band gaps at finite energy, making charge carrier confinement feasible. The gate-tunability of these phases allowed the demonstration of built-in Josephson junctions [18, 7, 9], superconducting quantum interference devices [19], Aharonov-Bohm rings [20] and superconducting channels [21]. However, well-controlled SETs and quantum dots - all of which require more complex gating structures - have not yet been demonstrated [22].
Here we demonstrate electrostatically controlled charge confinement and SET operation in tBLG based on an advanced gating structure. By using a first set of top gates, we create a narrow conducting channel in the tBLG by exploiting the single-particle band gaps (on the order of several tens of meV) separating the flat bands from the remote conduction and valence band. An additional set of narrow gates across the channel allows local control of the band edge profile and enables the formation of isolated charge islands, very similar as in Bernal-stacked bilayer QD devices [23, 24]. This makes it possible to demonstrate sequential tunneling through flat-band charge islands and using the associated SET as a platform to perform magnetic field spectroscopy measurements, providing information about the Fermi surfaces in the nearby reservoirs and their sensitive dependence on out-of-plane electric fields, confirmed by tight-binding calculations. In addition, our technology provides the basis for future studies of superconducting quantum point contacts, few electron quantum dots, Cooper pair splitters and related quantum devices in the tBLG.
Device operation
@@
Figure 2: Formation of a gate-defined SET. (a) Schematic illustration of the band edge profile along the gate-defined channel showing the formation of charge confinement between LG and RG by using the gap $\Delta_{+}$. (b) Large-range charge-stability diagram showing the conductance as a function of LG and RG measured at $V_{\mathrm{BG}}=3.5$ V and $V_{\mathrm{SG}}=3.5$ V (rhombus in Fig. 1d) with $V_{\mathrm{SD}}=100\,\mathrm{\mu V}$. (c) Line-cut showing the onset of Coulomb oscillations along LG for fixed $V_{\mathrm{RG}}=3.12$ V. (d) Zoom-in into the area marked by the black rectangle in panel (b) showing a set of diagonal Coulomb resonances indicating single-electron transport through a charge island located in between the LG and RG. (e) Schematic representation of the SET, with the charge island tunnel-coupled to the source (S) and drain (D) contacts and the corresponding gates highlighted. (f) Finite bias spectroscopy measurements for fixed $V_{\mathrm{RG}}=3.3301$ V [dashed line in panel (d)].
Finally, SET operation can be achieved by using both FGs to modulate the band edge profile along the conductive channel as illustrated in Fig. 2a. Fig. 2b shows a $V_{\mathrm{LG}}$-$V_{\mathrm{RG}}$ conductance map where we observe the transition from high conductance values (regime I, where the FGs are tuned such that we have flat-band carrier transport) to strongly suppressed conductance (regime II, where the FGs push the Fermi level into the gap $\Delta_{+}$). Focusing on a small gate voltage range applied to the LG and RG (black square in Fig. 2b and close-up in Fig. 2d), we observe well-defined conductance oscillations (see also Fig. 2c) that can be tuned equally well by the LG and RG (indicated by the slope of $\approx-1.08$ of the resonances in the $V_{\mathrm{LG}}$-$V_{\mathrm{RG}}$ plane). These periodic conductance oscillations (with a constant spacing of $\Delta V_{\mathrm{LG}}\approx 7.9\,$ mV), which we identify as Coulomb resonances, are a hallmark of the presence of tunneling barriers below both FGs, which consequently electrostatically define a charge island between the left and right FG. The occupation of this charge island, located in the flat bands, can be tuned by the capacitively coupled gates as illustrated in Fig. 2e and confirmed by finite bias spectroscopy measurements revealing characteristic Coulomb diamonds (Fig. 2f).
@@
Figure 3: Fully gate-tunable SET operation in the hole doped regime. (a) Schematic illustration showing the formation of a charge island between LG and RG by using the gap $\Delta_{-}$. (b) Charge-stability diagrams showing diagonal Coulomb resonances for different combinations of BG and SG voltages highlighting the flexibility of the device. Measurements are taken at (i) $V_{\mathrm{BG}}=-3\,\mathrm{V}$ and $V_{\mathrm{SG}}=-3.5\,\mathrm{V}$ at $V_{\mathrm{SD}}=150\,\mu\mathrm{V}$. (ii) $V_{\mathrm{BG}}=-4\,\mathrm{V}$ and $V_{\mathrm{SG}}=-3.23\,\mathrm{V}$ at $V_{\mathrm{SD}}=100\,\mu\mathrm{V}$. (iii) $V_{\mathrm{SG}}=-5\,\mathrm{V}$ and (slightly asymmetric) split gate voltages of $V_{\mathrm{SG1}}=-2.14\,\mathrm{V}$ and $V_{\mathrm{SG2}}=-1.94\,\mathrm{V}$ at $V_{\mathrm{SD}}=100\,\mu\mathrm{V}$ (circle in Fig. 1f) and (iv) $V_{\mathrm{BG}}=-6\,\mathrm{V}$ and $V_{\mathrm{SG}}=-1.7\,\mathrm{V}$ at $V_{\mathrm{SD}}=100\,\mu\mathrm{V}$. (c) Line-cut along the RG extracted from the left most panel in (b) for $V_{\mathrm{BG}}=-3\,\mathrm{V}$ at fixed $V_{\mathrm{LG}}=-4.925$ V showing periodic Coulomb oscillations with an approximate peak spacing of $\Delta V_{\mathrm{RG}}\approx 7.7$ mV. (d) Lower panel: Finite bias spectroscopy measurements as function of $V_{\mathrm{LG}}$ at fixed $V_{\mathrm{BG}}=-5\,\mathrm{V}$ and $V_{\mathrm{RG}}=-3.501$ V (see dashed line in central right panel in (b)). Upper panel: Zero bias Coulomb peak trace extracted from the Coulomb diamond measurement. Note the fact that complete current suppression can be achieved in this regime.
@@
The high tunability of the device is demonstrated by inverting all gate voltages and using the gap $\Delta_{-}$ to form local tunneling barriers essential for the SET operation, as illustrated in Fig. 3a. In Fig. 3b, we show corresponding $V_{\mathrm{LG}}$-$V_{\mathrm{RG}}$ conductance maps for four different $V_{\rm BG}$ values, where $V_{\rm SG}$ is set to ensure that the Fermi level is well within the gap $\Delta_{-}$ in the tBLG region below the SGs, and the LG and RG are fine-tuned to form tunneling barriers below the FGs (see also Supporting Section IV). In each case, we observe periodic Coulomb resonances (diagonal features), demonstrating the versatility of the device to create confinements using both single-particle band gaps of the tBLG electron system to realize a SET. Interestingly, we observe for all investigated regimes a nearly constant spacing of the Coulomb resonances of $\Delta V_{\mathrm{LG}}$ in the range of $(10.6\pm 3.2)$ mV (see e.g. conductance trace shown in Fig. 2c) and $\Delta V_{\mathrm{RG}}$ in the range of $(9.5\pm 2)$ mV, both consistent with the value observed for the electron-doped regime. This, in conjunction with the well-developed diagonal slopes with values in the range of $\approx-0.6$ to $\approx-1.1$ (Supporting Section II) and the possibility to completely suppress the conductance in certain gate voltage regimes, which results in the formation of well-developed Coulomb peaks (Fig. 3d), provides compelling evidence that the island is indeed formed consistently in the channel in between the two FGs. From the characteristic spacing of the Coulomb oscillations $\Delta V_{\mathrm{FG}}$ we can estimate the capacitance between the island and the FGs given by $C_{\mathrm{FG}}=e/\Delta V_{\mathrm{FG}}$ [25] resulting in $C_{\mathrm{FG}}$ in the range of $12$ to $22$ $\mathrm{aF}$. Interestingly, this value is more than a magnitude smaller than the self-capacitance extracted from the charging energy of the islands, which is consistently found to be in the range of $\approx 0.5$ meV (white dashed lines in Fig. 2f and lower panel in Fig. 3d), resulting in a self-capacitance of $C=e^{2}/E_{\mathrm{c}}\approx 320\,\mathrm{aF}$.
@@
Figure 5: Gate-tunable Fermi surface. (a) $B$-field derivative of the bulk conductance as a function of $V_{\mathrm{BG}}$ (bottom axis) or filling factor $\nu$ (top axis) and out-of-plane magnetic field. The dashed lines highlight the Landau levels emerging from CI at $\nu=-2$ (black) and the CNP at $\nu=0$ (yellow). (b) FFT of the $B$-field depnden quantum oscillations extracted for the different BG voltages (see Fig. 4d). (c) Quantum oscillation frequency $F_{B}$ obtained from tight-binding calculations for displacement field strengths of $D/\varepsilon_{0}=0$ V/nm (orange data) and $D/\varepsilon_{0}=0.48$ V/nm (dark red data). Note, that for converting $\nu$ to $V_{\mathrm{BG}}$ we neglect here stray fields from other surrounding gates. (d) Contour plots of the Fermi surface for different filling factors (i.e. energies) around the $K_{\mathrm{M}}$ points without (top row) and with (bottom row) an applied displacement field. (e) Schematic representation of the $D$-field induced energy shift. For more information see text.
@@
In conclusion, our measurements show that we can implement gate-defined SET in tBLG and use them as a platform to gain insights in band deformations of (near) magic angle tBLG in the presence of out-of-plane electric fields. Essentially, our extracted data fill the gap between the large-angle regime ($\theta\gg 1^{\circ}$), which is characterized by large bandwidths and comparatively little sensitivity to external electrical fields [40, 18] and the small-angle regime ($\theta\ll 1^{\circ}$), where the presence of electric fields causes the formation of topological helical networks [47, 48, 49]. An important observation is the absence of the correlated insulating state in Fig. 5b, while it is prominently present in the bulk transport data shown in Fig. 5a.
Given the recent theoretical and experimental study on the importance of long-ranged Coulomb interactions in the formation of correlated states in (near) magic-angle tBLG [50, 11, 42], we emphasize the flexibility of our device architecture to study electric field-induced screening of electronic interactions. In this respect, future experiments even closer to the magic angle could provide much-needed answers to a number of pressing questions related to correlated phenomena such as superconductivity and strange metal phases and their interplay with topology. In addition, the technology presented can enable the development of complex quantum devices and circuits based on tBLG, ranging from coupled quantum dots, Cooper-pair splitters, chains of Josephson junction-based SQUIDs to single-photon detector arrays.
@@
Author contributions C.S. and A.R. conceived the experiment. A.R. built the device and performed the measurements. A.F. performed the theoretical simulations. A.R. and A.A. performed the data analysis under the supervision of C.S.. E.I., K.H and L.B. provided assistance with the measurement setup. M.O. performed the ALD step. S.T. and F.L. performed the electron beam lithography. K.W. and T.T. supplied the hBN crystals. A.R., A.F., R.J.D., D.M.K., B.B. and C.S. discussed the experimental and theoretical data and wrote the manuscript.
Data availability The data supporting the findings of this study are available in a Zenodo repository under [insert DOI link here].